OpenStudy (anonymous):

in a 5.80E-2 M solution of a monoprotic acid HA, the acid is 23.9% dissociated. Calculate Ka for this acid.

OpenStudy (anonymous):

First you must find \[[H ^{+}] = \alpha \times cons of acid\] \[[H ^{+}] = 0.239 \times 0.058 = 0.0138\]

OpenStudy (anonymous):

Then \[[H ^{+}] =\sqrt{(Ka \times cons. of acid}\] \[0.0138 =\sqrt{(Ka \times 0.058}\] Ka = (0.0138)^2 / 0.058 = 3.31 x \[10^{-3}\]